OpenStudy (butterflydreamer):

Equilibrium Constant Calculation (simple one) - please check my working out :)

OpenStudy (photon336):

Where's the calculation?

OpenStudy (butterflydreamer):

oops sorry, internet cut out. Here it is: In a laboratory test of this equilibrium 6.00 mole of SO2 and 2.50 mole of O2 are added to 2.00L reaction vessel and allowed to come to equilibrium at 700 degrees celsius. If 4.00 mole of SO3 is present at equilibrium, what is the value of the equilibrium constant? \[2SO_2 (g) + O_2 (g) <=> 2SO_3 (g)\] So, i drew up an ICE table.  divide by 2 to obtain concentration:  So then the K constant: \[K_c = \frac{ [SO_3]^2 }{ [SO_2]^2 \times [O_2]}\] \[K_c = \frac{ [2.00]^2 }{ [1.00]^2 \times [0.25] } = 16\] I feel like i'm also meant to refer to the temperature (700 degrees celsius) but i'm not too sure? :/